# SeSE: A Structural Information-Guided Uncertainty Quantification Framework for Hallucination Detection in LLMs

### v2

@@ VIII-H Baselines Details
In-line Verbalized Confidence (IL-VC): In-line verbalized confidence (IL-VC) directly elicits the verbalized confidence about each claim $c$ in an in-line manner right after it is decomposed from the generations. Thus, we prompt LLMs with a long-form generation and instructions to give all the claims with corresponding confidence scores. The specific prompt that we use is as following:
@@
    The confidence score should represent your confidence in the claim, where a 1 is obvious facts and results like $1+1=2$. A 0 represents claims obviously incorrect or difficult for anyone to understand, such as “The Earth is the center of the universe” or “The exact population of a certain ordinary town”.
@@
P(True) variant. P(True) estimates the uncertainty of a claim by prompting LLMs to answer whether a claim is true or false, using the probability of the claim being true as the confidence score. Since probabilities cannot be obtained directly in long-form experiments, we improve this method by prompting the model to answer 10 times and estimating the probability by calculating the frequency of “true” responses. The specific prompts we use are as follows:
@@
SC-Prompt variant. This methods utilizes the consistency score of one claim across different samples. The consistency score is determined by prompting the same LLM.
Graph Centrality Metrics [19]. By integrating information theory with graph analysis, SeSE offers a principled measure of network structural disorder and establishes a computational paradigm for hierarchical abstraction. To validate SeSE’s advantages in quantifying intra-graph information, we compare it against various widely used graph-based analysis metrics. The specific definitions of these baseline metrics are detailed in Table VII, including Betweenness, Eigenvector, PageRank and Closeness.
TABLE VII: Benchmark graph centrality metrics with their formulas and explanations.
Metric Formula Brief Explanation
Eigenvector $C_{Eigv}(v)=\frac{1}{\lambda}\sum_{u\in N(v)}A_{vu}C_{Eigv}(u)$ Evaluates the influence of node $v$ based on the importance of its neighbors $N(v)$. $A_{vu}$ is adjacency matrix entry. $\lambda$ is the largest eigenvalue of the adjacency matrix.
TABLE VIII: The formulas and explanations of graph uncertainty metrics used in ablation study.
Metric Formula Brief Explanation
Eingvalue $U_{\text{Eigv}}=\sum_{k=1}^{m}\max(0,1-\lambda_{k})$ Quantifies graph uncertainty by summing positive differences between 1 and each Laplacian matrix eigenvalue $\lambda_{k}$, reflecting structural instability.
Degree $U_{\text{Deg}}=\frac{\text{trace}(|V|I-D)}{{|V|}^{2}}$ Measures graph uncertainty through normalized difference between ideal connectivity (scaled identity matrix $|V|I$) and actual connectivity (degree matrix $D$).
Eccentricity $U_{\text{Ecc}}=\frac{1}{m}\sum_{i=1}^{m}\left\|v_{i}-\bar{v}\right\|_{2}$ Captures graph uncertainty via average Euclidean distance between node embeddings $v_{i}$ and their mean $\bar{v}$, with larger distances indicating higher uncertainty.
Spectral Gap $U_{\text{Spec}}=1-\lambda_{2}$ $\lambda_{2}$ is the second smallest eigenvalue of the Laplacian matrix of $G$. Smaller $\lambda_{2}$ indicates weaker connections and higher uncertainty. $U_{\text{Spec}}$ inverts $\lambda_{2}$ to ensure that larger values reflect greater uncertainty.
  • $V$: The node set of graph $G$. AA: The adjacency matrix of graph $G$. $|V|$: The total number of nodes in $G$.

### v3

@@ G.2 Baselines in long-form experiments
  • In-line Verbalized Uncertainty (IL-VU): In-line verbalized uncertainty (IL-VU) directly elicits the verbalized confidence about each claim $c$ in an in-line manner right after it is decomposed from the generations. Thus, we prompt LLMs with a long-form generation and instructions to give all the claims with corresponding confidence scores. The specific prompt that we adapted from Mohri and Hashimoto [2024] is as following:
@@
    The confidence score should represent your confidence in the claim, where a 1 is obvious facts and results like $1+1=2$. A 0 represents claims obviously incorrect or difficult for anyone to understand, such as “The Earth is the center of the universe" or “The exact population of a certain ordinary town".
@@
P(True) variant. P(True) estimates the uncertainty of a claim by prompting LLMs to answer whether a claim is true or false, using the negative probability of the claim being true as the uncertainty score. Since probabilities cannot be obtained directly in long-form experiments, we improve this method by prompting the model to answer 10 times and estimating the probability by calculating the frequency of “true" responses. The specific prompt that we adapted from Kadavath et al. [2022] is as following:
@@
SC variant. This methods utilizes the consistency score of one claim across different samples, and the consistency score is determined by prompting the same LLM. The prompt used to evaluate consistency is described detailed in Appendix B.